40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jun 28 23:19:49 2018, johannes, Disha, Notes, Cavity, Gaussian beam mode for the new cavities New_Cavities.pdf
    Reply  Sun Jul 1 23:19:25 2018, johannes, Notes, Cavity, Gaussian beam mode for the new cavities 6x
       Reply  Tue Jul 3 11:58:25 2018, johannes, Notes, Cavity, Gaussian beam mode for the new cavities cavity_support_ring_gen2_15deg.PDFcavity_support_shelf.PDFshelf__contacts.pngcyl__contacts.png
          Reply  Wed Jul 11 22:38:54 2018, johannes, Notes, Cavity, Some assembly required 20180706_212353.jpg20180706_202052.jpg
             Reply  Wed Jul 11 23:27:38 2018, johannes, Notes, Cavity, new breadboard assembly 20180710_143617.jpg20180710_143629.jpg20180710_145112.jpg20180710_163221.jpg20180710_163803.jpg
                Reply  Wed Jul 11 23:51:18 2018, johannes, Notes, Cavity, first light 
                   Reply  Thu Jul 19 23:19:01 2018, johannes, Summary, Cavity, cryostat closed and pumping 20180719_114735.jpg20180717_164839.jpg
                      Reply  Sun Jul 22 00:52:10 2018, johannes, Summary, Cavity, transmission paths set up, mode-matching improved, post-baking ringdowns 6x
Message ID: 2103     Entry time: Wed Jul 11 23:51:18 2018     In reply to: 2102     Reply to this: 2104
Author: johannes 
Type: Notes 
Category: Cavity 
Subject: first light 

[Disha, Johannes]

Today we aligned the incident beams to the cavities and locked both lasers for the first time. As a reminder, this is the procedure we're following:

  1. Estimate BB lens positions
  2. Align beams into cavities
  3. Lock lasers
  4. With lasers locked, tune the placement of the BB lenses on the transmission side for parallel, level beams with 5" separation and 4" beam height above the optical table to hit the viewports (BB is in nominal suspended position for this, but propped up on posts, with springs not attached)
  5. Turn the BB around
  6. Rinse & repeat to tweak the input lens placement as well
  7. Turn breadboard again
  8. Optimize input mode-matching
  9. Close cavity cans
  10. Suspend breadboard
  11. BB balancing

For now the input mode-matching is not yet optimized, just placed approximately following some initial calculations. Reflected light drop was of order 60-70% and the lasers locked just fine with plenty of light being transmitted (easily visible on card = good for aligning lenses). Work to be continued tomorrow.

Some good news: In air the beat frequency between the locked lasers was 446 MHz, which is on the lower end of expected natural resonance differences. With the old cavities, the cryogenic beat frequency was within 10% of that at room temperature with no thermal control. The FSR of the new cavities is 4 GHz, so it looks we're quite lucky: we're in the top 25 percent of favorable beat frequencies, well within the bandwidth of our photodetectors, and can see the signal on AG4395's. Hopefully this is true at cryo as well.

 

ELOG V3.1.3-